When Naomi was born, her parents invested $1000 in a fixed rate savings account at a rate of 7% compounded monthly. Naomi will receive when she turns 18 to help with her college expenses. What amount of money will she receive. Round your answer to the nearest dollar.
Answer by Boreal(15235)

A=Ao(1+(.07/12))^216, where Ao=$1000 and 216 comes from 18 years*12 months/year, the number of compoundings.
calculator, start with 1+(.07/12), raise it to the 216 power, and then multiply by 1000.
$3512.54 or $3513.
Check: rule of 72 for doubling is 72/interest rate =years needed or 72/7=10.3
Not quite 2 doublings, which means not quite $3000. This is a reasonable answer.